OpenStudy (anonymous):

Write three fractions that are equivalent to 6/9 Show your work

OpenStudy (kymber):

Divide or multiply both the numerator & denominator by the same number

Parth (parthkohli):

If we multiply the same number to both the numerator and denominator, we get an equivalent fraction. Can you do it now? ;)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

but i need the answer

Parth (parthkohli):

Okay 6/9 = 6*2/9*2 = 12/18 6/9 = 6*3/9*3 = 18/27 6/9 = 6*4/9*4 = 24/36 Okay?

OpenStudy (anonymous):

thank you